OpenStudy (anonymous):

17/21 - (2/7 + 1/7 ) will become a fan a. 2/3 b. 1 5/21 c. 2/21 d. 8/21

OpenStudy (gorv):

\[\frac{ 17 }{ 21 }-(\frac{ 2 }{ 7 }+\frac{ 1 }{ 7 })\]

OpenStudy (gorv):

last two number have same denominator so add njumerator directly ...tell me what u get??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3

OpenStudy (gorv):

so 3 /7

OpenStudy (gorv):

\[\frac{ 17 }{ 21 }-\frac{ 3 }{ 7 }\]

OpenStudy (gorv):

here both number have different denominator so will take LCM

OpenStudy (gorv):

u know abt lcm??

OpenStudy (anonymous):

yes

OpenStudy (gorv):

what will be lcm of 7&21

OpenStudy (anonymous):

8/21

OpenStudy (anonymous):

would that be the answer

OpenStudy (anonymous):

?

OpenStudy (gorv):

yeahhh ..buddy u r over it ...very good

OpenStudy (anonymous):

can you help me with 4/7 / 1.5

OpenStudy (gorv):

ans is right

OpenStudy (anonymous):

thank you i already did the other one